Problem 9. Let n be the number of drive-in movie sites in the United States at t years since 2009. The function n 49t +381 models the situation well for the period 2009-2014 (Source: National Association of Theater Owners) a) Rewrite the equation n- 4.9t+381 using the function name f b) Find 13). What does it mean in this situation? c) Find f0). What does it mean in this situation?
Answer #1

(a)

The given function is rewritten using the function name f as follows:

n_{i}=f(t_{i})=-4.9t_{i}+381,

for i = 0,1,2,...

where ti refers to ith year after 2009.

(b)

For i = 3, we have:

n_{3}=f(t_{3})=-4.9t_{3}+381=-(4 imes 3)+381=369

In this situation, it means the number of drive-in movie sites in United States in the year 2012 = 369

(c)

For i = 0, we have:

no f(to) _ _4.9to+ 381--(4 x 0) + 381-381

In this situation, it means the number of drive-in movie sites in United States in the year 2009 = 381..